answersLogoWhite

0

well since the coins have two sides,there is a 50% chance of it landing on heads

User Avatar

Wiki User

∙ 14y ago

What else can I help you with?

Related Questions

Two coins are tossed What is the probability of both coins are heads?

The probability that both coins are heads is the probability of one coin landing heads multiplied by the probability of the second coin landing heads: (.5) * (.5) = .25 or (1/2) * (1/2) = 1/4


What is the probability of flipping three coins and all of them land up on heads?

The probability of the first coin landing heads is half (or 1/2). Similarly, the probability of the second and third coins landing heads are also 1/2 in each case. Therefore, the probability of having three heads is: (1/2)(1/2)(1/2) = (1/8)


What is the probability of flipping two coins and one landing on heads and one landing on tails?

It is 1/2.


2 coins are flipped. What is the probability of landing one heads and one tails?

It is 1/2.


What is the probability of seven coins tossed simultaneously and all of them landing on heads?

It is 1/2^7 = 1/128.


A fair nickel and a fair dice are tossed once outcomes match when both coins land on heads or both coins land on tails What is the theoretical probability of a match?

The probability of 2 coins both landing on heads or both landing on tails is 1/2 because there are 4 possible outcomes. Head, head. Head, tails. Tails, tails. Tails, heads. Tails, heads is different from heads, tails for reasons I am unsure of.


What is the probability of getting 2 heads when 2 coins are tossed?

The probability of tossing two heads in two coins is 0.25.


What is the probability of landing on heads twice?

The probability of landing on heads each time a fair coin is flipped, is 1/2.Assuming that the question was supposed to be:"What is the probability of landing on heads twice in a row?"To calculate compound probabilities like this, we first have to work out the probability of landing on heads each time, and then multiply these two probabilities to get a compound probability.1/2 x 1/2 = 1/4So the probability of landing on heads twice in a row = 1/4 (for a fair coin)


What is the probability of landing on an even number when tossing an 8 sided die and tossing a coin and getting heads?

1/2 * 1/2 = 1/4 1/2= probability of landing an even number 1/2 = probability of landing a heads


Three fair coins are tossed determine the probability of the following event exactly two coins land on heads?

The sample space is 23 or 8; which can be listed out as: HHH, HHT, HTH, HTT, THH, THT, TTH, TTT. There are 2 of the 8 that have exactly 2 heads; so the probability of exactly two coins landing on heads is 2/8 or 1/4.


What is the probability of a coin landing on heads?

1 in 2, or 0.5


What is the probability of landing on heads 6 times?

The probability is 6 in 12, or 1 in 2.

Trending Questions
Is it true the special rule of addition is the events must be mutually exclusive? How many tens are in 15000000? A sample taken without bias? How many permutations of letters in word swimming? What is the top 1 percent income in us? If a rifleman hits target 80 percent of the time If he fires four shots what is the probability he will hit the target exactly twice and what is the probability he will hit the target at least twice? A box contains four balls two red and two yellow balls we extract two balswhat is the probability of extracting two red balls? Which cursor attribute provides you with the number of records affected by the last DML statement? How many 3 digit numbers have the sum of their digits equal to 5? Why are sample used in statistics? Which step in composite risk management is adtp-tl used? What is initial temperature mean? What is the probability drawing an ace followed by a king? The parameter of interest when conducting a test of significance for a proportion is? A boy has 4 red and 4 yellow and 5 green marbles In how many ways can the boy arrange the marbles in a line if all marbles of the same color are indistinguishable? The trapezoidal rule for integration gives exact result when the integrand is a polynomial of degree? What is the different between status and rank? What is the off chance of someone being a silent sleeper? In both elective and merit selection systems a person from which group is the most likely to be selected? Is genotype quantitative or qualitative?